# Can LLMs Reason Structurally? An Evaluation via the Lens of Data Structures

### v2

@@ Abstract
As large language models (LLMs) take on increasingly complex tasks, understanding their algorithmic reasoning abilities has become essential. However, existing evaluations focus on distinct and isolated tasks. We propose a unified diagnostic lens: structural reasoning—understanding and manipulating relationships like order, hierarchy, and connectivity. We introduce DSR-Bench, the first benchmark to systematically evaluate LLM structural reasoning through canonical data structures, which serve as interpretable, algorithmically meaningful abstractions. DSR-Bench spans 20 data structures, 35 operations, and 4,140 synthetically generated problem instances with minimal contamination. The benchmark’s hierarchical design pinpoints specific failure modes, while its fully automated evaluation ensures objective and consistent assessment. Benchmarking ten state-of-the-art LLMs reveals critical limitations: the top-performing model scores only 0.498 out of 1 on challenging instances. Three additional evaluation suites reveal further weaknesses: models perform poorly on spatial data and natural language scenarios, and fail to reason over their own generated code. DSR-Bench offers a principled diagnostic tool for structural reasoning, helping expose reasoning bottlenecks and guide the development of more capable and reliable LLMs.
@@ 1 Introduction
Figure 1: Overview of DSR-Bench’s main suite with six data structure categories capturing distinct relationships, plus the challenge subset. Three specialized suites that holistically evaluate structural reasoning under different settings: spatial (multi-dimensional data), natural (realistic natural language scenarios), and code (code generation).
As large language models (LLMs) tackle increasingly complex real-world challenges, strengthening their algorithmic reasoning abilities is essential for interpretability, safety, and efficiency (Eberle et al., 2025). Yet scaling alone is reaching its limits: models face data scarcity and diminishing returns (Villalobos et al., 2024). A complementary path is to study the algorithms LLMs learn and employ, which can reveal their internal representations and inspire algorithm-centric architectures that embed reasoning capabilities directly into model design (Bounsi et al., 2024; Eberle et al., 2025).
@@
The call to prioritize algorithmic reasoning has intensified as the field shifts toward evaluating general reasoning without external tools. While LLMs can solve many tasks through code generation and execution, recent initiatives like Gemini-Deep-Think’s and OpenAI’s IMO competitions (Luong & Lockhart, 2025; Wei, 2025) explicitly prohibit coding and proof assistance. These efforts highlight a growing emphasis on end-to-end reasoning as a critical step toward artificial general intelligence.
Despite rapid progress, existing evaluations of algorithmic reasoning remain fragmented, each probing only a single facet, such as arithmetic operations (Zhou et al., 2022; 2024; Lee et al., 2024), simulation of textbook algorithms (Markeeva et al., 2024), or domain-specific tasks like graph problems (Wang et al., 2023a; Fatemi et al., 2024). Meanwhile, interpretability research, such as mechanistic interpretability (Olah et al., 2020; Wang et al., 2023b), examines internal model representations rather than reasoning over algorithmic structures. What is missing is a unified framework that bridges these threads, offering interpretable diagnostics of algorithmic reasoning across diverse problem types.
@@
To address this gap, we propose to evaluate algorithmic reasoning through the lens of data structures—fundamental abstractions that organize information and govern algorithmic operations. Data structures provide an ideal evaluation framework because they are both systematic (covering diverse relationship types) and interpretable (offering clear diagnostic insights). They span the spectrum of relationships underlying algorithmic thinking: arrays represent sequences, stacks and queues capture temporal ordering, trees encode hierarchies, and graphs represent complex networks. This diversity enables us to holistically assess whether models can construct, maintain, and manipulate different types of relationships—the core capability we term structural reasoning.
@@
Structural reasoning underlies many real-world applications. Trip planning requires interpreting maps as graphs and managing priorities with queues, while supply chain optimization depends on hierarchical resource allocation and temporal sequencing. In safety-critical domains such as robotics (Sado et al., 2023) and healthcare (Sadeghi et al., 2024), failures in structural reasoning abilities can have serious consequences, making systematic evaluation crucial before deployment.
Contributions
@@
We introduce DSR-Bench, the first benchmark to systematically isolate and evaluate LLM structural reasoning through data structure manipulation tasks. DSR-Bench comprises 4,140 problem instances spanning 20 data structures (grouped into six relationship categories; see Figure˜1) and 35 operations across three difficulty levels (short, medium, and long) and five evaluation suites (main, challenge, spatial, natural, and code). Each task provides interpretable, automatically verifiable assessments of whether models can construct, maintain, and reason about relationships. The highlights and main takeaways of DSR-Bench are:
@@
  • Hierarchical and diagnostic design. Simpler tasks (e.g., queue operations) serve as prerequisites for complex ones (e.g., graph traversal). This enables fine-grained localization of failure modes.
  • Deterministic, contamination-resistant evaluation. Synthetic data minimizes contamination risk. The evaluation is deterministic and fully automated, supporting objective and fair assessment.
@@
  • Five evaluation suites covering diverse reasoning settings.
@@
      main: Canonical data structure tasks reveal core reasoning gaps—e.g., instruction-tuned models struggle with multi-attribute and multi-hop reasoning, while reasoning models fail to override memory for non-standard tasks with user-defined constraints. Simpler prompts consistently help, but CoT requires careful design.
@@
      challenge: Hybrid and compositional structures remain difficult, with the best model scoring only 0.498 out of 1, revealing reasoning limitations of frontier models.
@@
      spatial: On multi-dimensional data, the common format found in real-world applications, models struggle as dimensionality increases or when data follows non-uniform distributions.
@@
      natural: In realistic natural language scenarios, models struggle to extract structure and navigate ambiguity, exposing a gap to real-world deployment.
@@
      code: An auxiliary probe contrasting reasoning with and without tool use. Models rarely benefit from reasoning over their own generated code; external interpreters help on familiar tasks but consistently fail on non-standard or natural-language variants.
  • Comprehensive empirical analysis. We evaluate ten state-of-the-art LLMs and analyze prompting strategies, distribution shifts, and qualitative error patterns (e.g., implicit priors, instruction-following failures). These findings provide actionable insights for developing architectures and training methods that better capture algorithmic reasoning.
To welcome community engagement and collaboration, we release all code at https://github.com/dransyhe/DSR-Bench and data at https://huggingface.co/collections/vitercik-lab/dsr-bench-6826381f6297ff1499134163 for full reproducibility.
2 Related work
Algorithmic reasoning with LLMs
Prior works target distinct but isolated facets of algorithmic reasoning, often limited in scope: arithmetic tasks and length generalization in transformers (Zhou et al., 2022; 2024; Lee et al., 2024), or specific graph problems like cycle detection and connectivity (Wang et al., 2023a; Fatemi et al., 2024). CLRS-Text (Markeeva et al., 2024) evaluates whether LLMs can simulate 30 classical algorithms (Veličković et al., 2022; Cormen et al., 2009). We take a holistic and foundational approach using data structures, which serve as interpretable building blocks of algorithms and support fine-grained diagnosis of specific reasoning failures to inform future research.
Reasoning benchmarks
Existing LLM reasoning benchmarks are predominantly high-level and domain-specific, targeting math (Cobbe et al., 2021; Liu et al., 2024a; b), STEM (Hendrycks et al., 2021), and logic puzzles (White et al., 2025; Giadikiaroglou et al., 2024). These often require complex responses with intertwined reasoning steps, relying on subjective human or LLM-based evaluation (Chiang et al., 2024; Feuer et al., 2024; Ye et al., 2024). We focus on structural reasoning, an implicit requirement underlying problem-solving across domains. By using data structures as clear abstractions of these relationships, we isolate structural reasoning from domain-specific complexities.
Coding benchmarks
Coding benchmarks evaluate how well LLMs write syntactically correct code or function as coding agents (Chen et al., 2021; Zheng et al., 2023; Jimenez et al., 2024; Jain et al., 2025; White et al., 2025; Aider-AI, 2025), typically requiring external interpreters for verification. While useful, these benchmarks conflate reasoning with tool execution and are limited to domains where coding applies. In contrast, we target general reasoning independent of external tools, reflecting the broader goal of assessing progress toward general intelligence. To bridge these perspectives, we include the code suite to probe whether code generation aids structural reasoning.
@@ 3 DSR-Bench: the Data Structure Reasoning Benchmark
Figure 2: Left: Scores of ten models on DSR-Bench-main, averaged across three runs. Right: Radar chart showing scores of top-performing models across six data structure categories. We note DSR-Bench also includes a challenge suite, where the best model scores only 0.498.
In DSR-Bench, we propose a taxonomy of data structures grouped into six categories based on the types of data relationships they encode: Linear, Temporal, Associative, Hierarchical, Network, and Hybrid. This taxonomy captures the diversity of relational patterns found in real-world data, enabling systematic evaluation across distinct reasoning challenges.
@@ 3.1 Tasks
DSR-Bench includes the following data structures and tasks, further detailed in Section˜A.2.
@@
  • Linear (Sequential): This category includes array and its operations: access, insert, delete, reverse, and search. Linear structures introduce ordered relationships, enabling reasoning about position, sequence, and iteration. They serve as a foundation for more complex data abstractions.
  • Temporal (Time-based ordering): Temporal structures include stack, queue, priority queue, and skip list, which operate under last-in-first-out, first-in-first-out, priority-based, or probabilistic rules. Their operations are compound actions of insertions and deletions. Temporal reasoning is essential in systems that require ordered execution over time, such as event queues and schedulers.
@@
  • Associative (Key-value mapping): We evaluate hashmap, trie, and suffix tree through tasks focused on construction and compound insertion-deletion operations. Tasks test the ability to handle key-based insertions, retrievals, and pattern matching in nested or hashed structures. These skills are core for efficient lookup and structured access in systems such as databases.
  • Hierarchical (Tree-like): This group includes binary search tree (BST), heap, red-black (RB) tree, and B+ tree, with tasks involving traversals and compound operations. These structures are common in file systems and database indexing. Tasks assess whether the model can maintain structural invariants, perform updates efficiently, and simulate recursive behavior.
@@
  • Network (Connectivity and group membership): This category includes Graph and Disjoint Set Union (DSU) tasks. Graph tasks cover breadth-first and depth-first traversal, testing the ability to reason about connectivity and explore many-to-many relationships, as commonly found in social networks. DSU tasks evaluate union–find operations, which are essential for network connectivity analysis and clustering algorithms.
@@
  • Hybrid (Combined relationships): Real-world systems often rely on hybrid data structures that combine multiple forms of structural reasoning. This category includes LRU cache, which integrates temporal and memory management; bloom filter, for probabilistic set membership; and Directed Acyclic Word Graph (DAWG), a trie-like hierarchical graph. Tasks test whether models can compose different principles and generalize beyond individual structures.
Hierarchical organization
Tasks are organized by increasing complexity for fine-grained diagnosis of reasoning failures. For example, success on list but failure on hashmap isolates the issue to multi-attribute access rather than linear sequencing. Similarly, since tree traversal depends on queue or stack operations, testing these primitives becomes a prerequisite.
Operations
@@
We design a diverse set of operation tasks for each data structure (full list in Table˜10), spanning three categories: construction, inspection (e.g., access, traversal), and manipulation (e.g., insert, delete). Beyond these atomic operations, DSR-Bench includes compound operations: sequences of manipulations (e.g., insert, insert, delete, …) designed to test whether models can maintain structural integrity across multiple steps.
@@
We assign tasks to three difficulty levels based on input length: short (5–10), medium (11–20), and long (21–30), to assess length generalization. For atomic operations, length refers to the number of input elements (e.g., list items, tree nodes). For compound operations, tasks begin with an empty structure, and length is defined by the number of sequential operations.
Evaluation suites
Beyond the main suite, we curated a challenge subset targeting particularly complex structures to stress-test advanced reasoning capabilities. To more comprehensively evaluate structural reasoning beyond canonical data structures, we introduce three additional suites: spatial, which evaluates reasoning on multi-dimensional data; natural, which tests reasoning when tasks are embedded in realistic natural language scenarios; and code, which assesses whether models can leverage code generation for structural reasoning. Each suite examines how well structural reasoning capabilities transfer to scenarios critical for real-world deployment.

### v3

# Can LLMs Reason Structurally? Benchmarking via the lens of Data Structures

@@ Abstract
Large language models (LLMs) are deployed on increasingly complex tasks that require multi-step decision-making. Understanding their algorithmic reasoning abilities is therefore crucial. However, we lack a diagnostic benchmark for evaluating this capability. We propose data structures as a principled lens: as fundamental building blocks of algorithms, they naturally probe structural reasoning—the ability to understand and manipulate relationships such as order, hierarchy, and connectivity that underpin algorithmic reasoning. We introduce DSR-Bench, spanning 20 data structures, 35 operations, and 4,140 problem instances. DSR-Bench features hierarchical task organization, fully automated generation and evaluation, and fine-grained diagnostics. Evaluating 13 state-of-the-art LLMs reveals critical limitations: the top-performing model achieves only 0.46/1 on challenging instances. Three auxiliary probes targeting more realistic usages expose further weaknesses: models perform poorly on spatial data and context-rich scenarios, and they struggle to reason over their own code.
Machine Learning, ICML
@@ 1 Introduction
As large language models (LLMs) tackle increasingly complex real-world tasks, it is essential to understand their algorithmic reasoning abilities (Eberle et al., 2025)—that is, their ability to understand, learn, and execute algorithmic operations. By studying the algorithmic primitives that LLMs learn and employ, we can analyze their reasoning processes and how they compose these primitives in multi-step problem-solving. Such insights may also inspire algorithm-centric model design as an alternative to the scaling paradigm (Bounsi et al., 2024; Eberle et al., 2025).
@@
In our analysis of algorithmic reasoning, we explicitly focus on LLMs’ inherent reasoning abilities, isolated from external tools. This emphasis is motivated by recent initiatives like Gemini-Deep-Think’s and OpenAI’s participation in IMO competitions (Luong and Lockhart, 2025; Wei, 2025), which strictly prohibit the use of code and proof assistance, emphasizing end-to-end reasoning as a step toward artificial general intelligence.
Limitations of existing reasoning and coding benchmarks.
Most reasoning benchmarks focus on high-level, domain-specific tasks such as mathematics (Liu et al., 2024b, a), data analysis (White et al., 2025), or STEM questions (Hendrycks et al., 2021). These tasks often require complex responses with intertwined reasoning skills, making it difficult to isolate algorithmic primitives. Coding benchmarks such as LiveCodeBench (Jain et al., 2025), SWEBench (Jimenez et al., 2024), and Aider (Aider-AI, 2025) assess code generation, which risks contamination due to the abundance of online resources. Furthermore, coding offloads essential reasoning processes to external interpreters, obscuring the inherent reasoning abilities we aim to study. We therefore need a fundamental approach to isolate algorithmic reasoning from domain-specific complexities and evaluate it without using external tools.
Limitations of existing algorithmic benchmarks.
Despite growing research on algorithmic reasoning, existing benchmarks fall short in enabling systematic, granular assessment. CLRS-Text (Markeeva et al., 2024) evaluates the simulation of 30 classical algorithms (Cormen et al., 2009), but empirical differences across algorithms (e.g., success on bubble sort versus failure on insertion sort) provide limited insights into why these differences arise or which underlying operations are driving them. Additionally, its prompts are direct textual translations of the CLRS-30 Benchmark (Veličković et al., 2022), originally designed for non-textual models such as graph neural networks: they use code-like syntax and provide no natural-language task description beyond the algorithm name. Existing algorithmic benchmarks with natural language prompts are primarily limited to graphs (Wang et al., 2023; Fatemi et al., 2024). Therefore, we need a comprehensive natural-language benchmark over diverse structures that enables fine-grained diagnosis, beyond end-to-end algorithmic simulation.
Key challenge.
The central challenge of curating a diagnostic benchmark that addresses these limitations is defining algorithmic primitives (Eberle et al., 2025) that enable decomposition and failure localization. We need a framework that (i) provides a principled dictionary of fundamentals, (ii) supports the composition of basic skills for complex reasoning, and (iii) enables automated, deterministic, and unambiguous evaluation. This is difficult because these requirements are in tension: tasks should be atomic to prevent heuristic or shortcut solutions and enable precise error attribution to satisfy (i), yet expressive enough to cover a wide range of complex tasks for (ii). Moreover, satisfying (iii) is challenging because deterministic, unambiguous evaluation is hard in natural language, where errors can be confounded by unspecified tie-breaking or prompt ambiguity.
Why data structures?
@@
We propose data structure tasks as a well-suited testbed for algorithmic reasoning. Data structures are fundamental to algorithms, interpretable in their operations, and deterministic in their outputs. They also abstract the key relationships underlying algorithmic reasoning: arrays represent sequences, trees encode hierarchies, and graphs capture complex networks. This diversity enables us to holistically assess whether models can construct, manipulate, and compose different relationships—the core capability we term structural reasoning, which we use as a framework for probing algorithmic reasoning.
@@
More importantly, structural reasoning is a fundamental requirement for many real-world applications. For example, trip planning requires interpreting maps as graphs and managing scheduling priorities with queues, while supply chain optimization depends on hierarchical resource allocation and temporal sequencing. Understanding how well LLMs reason structurally is thus essential for their safe deployment in critical domains such as robotics (Sado et al., 2023) and healthcare (Sadeghi et al., 2024).
Figure 1: High-level overview of DSR-Bench-main with six data structure categories capturing distinct relationships, plus the challenge subset. Three specialized probes that holistically evaluate structural reasoning under different settings: spatial (multi-dimensional data), realistic (realistic context-rich scenarios), and code (code generation). See Table˜5 for full list of tasks.
Our contribution.
@@
We introduce DSR-Bench, the first benchmark for systematically evaluating structural reasoning in LLMs through data structure tasks. It comprises 4,140 problem instances spanning 20 data structures, 35 operations, three difficulty levels (short, medium, and long), and five evaluation components (main, challenge, spatial, realistic, and code). Data structures are grouped into six categories (see Figure˜1), and each task enables interpretable, automatically verifiable assessments of whether models can reason over a specific relationship. The highlights and main takeaways of DSR-Bench are:
@@
  • Hierarchical task organization. Tasks are organized in increasing difficulty for structural reasoning, where simpler tasks (e.g., queues) serve as prerequisites for complex ones (e.g., breadth-first traversal on trees). This design enables fine-grained localization of failure modes.
  • Deterministic, automated evaluation. All data is generated synthetically, enabling efficiency while minimizing contamination risk. Evaluation is fully automated with unambiguous ground truth, avoiding human or LLM-based judging for fair assessment.
  • Interpretable, diagnostic insights. Data structures have well-defined semantics and easily computable ground truth. This allows for direct comparisons of model reasoning traces and interpretable analyses of how models reason and where they fail, mitigating leaderboard-driven reward hacking (Amodei et al., 2016).
@@
  • Five evaluation components covering diverse reasoning settings, offering practical takeaways.
@@
      main: Canonical data structure tasks. Empirical analysis reveals several gaps, e.g., instruction-tuned models struggle with multi-attribute and multi-hop reasoning, while reasoning models fail to adapt to user-defined constraints. Simpler prompts consistently help, though CoT requires careful design.
@@
      challenge: Complex structures (e.g., hybrid and compositional) with longer inputs. The best-performing model scores only 0.46 out of 1, revealing limitations in frontier models.
@@
      spatial: High-dimensional data structural tasks common in real-world applications. Results show that models degrade as dimensionality increases and struggle with non-uniform distributions.
@@
      realistic: Data structures embedded in realistic scenarios (e.g., clinic appointments). We find that models struggle to extract structure and navigate language ambiguity, exposing a significant gap between formal reasoning and real-world deployment.
@@
      code: A supplementary probe evaluating structural reasoning with and without code generation. Though our focus is on LLMs’ inherent reasoning, we include this probe to assess whether code provides additional benefit. We observe that models rarely benefit from reasoning over self-generated code. External interpreters help with familiar tasks but fail on non-standard or realistic variants.
  • Comprehensive empirical analysis. We evaluate 13 state-of-the-art LLMs spanning open- and closed-source, instruction-tuned and reasoning models. Our ablations examine prompting strategies, distribution shifts, and qualitative error patterns (such as testing implicit priors and instruction-following failures). Detailed empirical analysis is presented in Section˜4. These findings offer actionable insights for designing architectures and training strategies to improve algorithmic reasoning in LLMs.
  • Open-source. We release all code at https://github.com/dransyhe/DSR-Bench and datasets at https://huggingface.co/collections/vitercik-lab/dsr-bench for community engagement.
2 Additional related work
We review the limitations of related benchmarks in Section˜1, with an extended discussion for LLM reasoning and coding benchmarks in Appendix˜A.
Table 1: Comparison with prior algorithmic benchmarks.
Benchmark Coverage Hierarchical Atomic Expressive Natural language Auxiliary probes
CLRS-Text 30 algorithms
NLGraph Graphs only
GraphQA Graphs only
DSR-Bench 20 DS, 6 categories
In Table˜1, we summarize how DSR-Bench differs from prior algorithmic benchmarks, where “atomic” and “expressive” correspond to requirements (i) and (ii) discussed in Section˜1’s Key Challenge. We further elaborate on task-level distinctions. NLGraph (Wang et al., 2023) and GraphQA (Fatemi et al., 2024) focus on graph tasks (e.g., connectivity, cycle detection). While DSR-Bench also has graph tasks (breadth-first search (BFS), depth-first search (DFS)), these tasks do not appear in those benchmarks and are integral parts of our hierarchical design, serving as preliminaries for harder tasks (e.g., directed acyclic word graph). CLRS-Text (Markeeva et al., 2024) evaluates end-to-end simulation of 30 classical algorithms from Cormen et al. (2009) (e.g., sorting, greedy algorithms), with only BFS and DFS overlapping with ours. In contrast, DSR-Bench targets data structures as core building blocks underlying these algorithms, enabling finer-grained diagnosis and more interpretable, actionable insights.
@@ 3 DSR-Bench: the Data Structure Reasoning Benchmark
We detail the design of DSR-Bench. In Section˜3.1, we describe the task composition, which spans six relationship categories. We then present the task organization and five evaluation components (Section˜3.2), prompt design (Section˜3.3), and data generation and evaluation (Section˜3.4).
Figure 2: Left: Scores of thirteen models on DSR-Bench-main, averaged across three runs. Right: Radar chart showing scores of top-performing models across six relationship categories. Note that the best model on DSR-Bench-challenge scores only 0.46.
@@ 3.1 Tasks
We propose six relationship categories fundamental to algorithm reasoning: Linear, Temporal, Associative, Hierarchical, Network, and Hybrid. We design data structure tasks to cover each category; see Appendix˜B for their details.
@@
  • Linear (Sequential): This category includes array and its operations (access, insert, delete, reverse, search). It captures basic ordering of elements, which forms the basis of more complex structures.
  • Temporal (Time-based ordering): Temporal includes stack, queue, LRU cache, and priority queue, imposing additional constraints on top of linear structures. Our tasks include compound sequences of insertions and deletions. Temporal structure is often used in time-sensitive applications, including scheduling.
@@
  • Associative (Key-value mapping): This category includes hashmap, trie, suffix tree, and skip list. We evaluate the ability to reason about associative relationships through construction and compound operations. Associative structures are essential for efficient lookup and access in systems such as databases.
  • Hierarchical (Tree-like): This group includes binary search tree (BST), heap, red-black (RB) tree, and B+ tree, testing traversal and compound operations. These structures require data maintenance over multiple level sets and are common in file systems and databases.
@@
  • Network (Connectivity and group membership): This category includes Graph and Disjoint Set Union (DSU) tasks. Graph tasks include traversals, testing reasoning over many-to-many relationships, as in social networks. DSU tasks test union–find operations used in connectivity and clustering algorithms.
@@
  • Hybrid (Combined relationships): Real-world systems often require a combination of different structural reasoning skills. This category includes Bloom filter (probabilistic set membership) and Directed Acyclic Word Graph (DAWG) (trie-like hierarchical graph). These tasks test whether models can compose and generalize beyond individual structures.
3.2 Benchmark design
Hierarchical task organization
A key strength of DSR-Bench is that it organizes tasks by increasing complexity, where simpler tasks are prerequisites for more complex ones. For example, performing a DFS (Depth-First Search) on a graph requires stack operations to maintain the frontier, enabling fine-grained diagnoses that isolate where reasoning begins to fail.
Operation types
@@ 3.2 Benchmark design
For each data structure, DSR-Bench provides a diverse set of operation tasks (see the full list in Table˜5), spanning three categories: construction, inspection (e.g., access, traversal), and manipulation (e.g., insert, delete). Beyond these atomic operations, it also includes compound operations, which are sequences of operations (e.g., [insert, insert, delete, …]) designed to test whether models can perform multi-step structural reasoning.
@@
Tasks are assigned to three difficulty levels based on input length: short (5–10), medium (11–20), and long (21–30), to assess length generalization. For atomic operations, length refers to the number of input elements (e.g., tree nodes). For compound operations, length is defined by the number of sequential operations.
Evaluation components
Beyond the main suite, we curate a challenge subset with particularly complex structures to stress-test advanced reasoning. We further introduce three targeted probes to evaluate structural reasoning in scenarios applicable to real-world deployment: spatial, evaluating on high-dimensional data; realistic, embedding data structure tasks in realistic scenarios with context-rich language; and code, assessing whether models can leverage code generation to aid structural reasoning.